find the integral surface of the equation (x-y)y^2*p + (y-x)x^2*q = (x^2+y^2)z through the curve xz = a^3 and y = 0


1
Expert's answer
2021-04-29T16:59:02-0400

Auxiliary equations

"\frac{dx}{(x-y)y^2}=\frac{dy}{(y-x)x^2}=\frac{dz}{(x^2+y^2)z}"

Taking dx and dy

"\frac{dy}{dx}=-\frac{x^2}{y^2}"

"y^2dy=-x^2dx"

"C_1=x^3+y^3"

"\frac{dx}{(x-y)y^2}=\frac{dz}{(x^2+y^2)z}"

"\frac{(x^2+y^2)dx}{(x-y)y^2}=\frac{dz}{z}"

"lnz=\frac{1}{y^2}F(x,y)+C_2"

Here, y cannot be 0 in this equation that is why there is no intersection with curve

xz=a3, y=0